OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

okay give me a sec

OpenStudy (anonymous):

So I got 3/1

OpenStudy (anonymous):

So, you got \(x=3\) ?

OpenStudy (anonymous):

I'd recheck that work.

OpenStudy (phi):

sith made a typo the problem is 4x+2(1-2x)= -1

OpenStudy (anonymous):

yes i guess but that is not one of my anwer choices

OpenStudy (anonymous):

okay let me work it out

OpenStudy (anonymous):

Ah, thanks for pointing it out @phi. In any case, it's still not right.

OpenStudy (anonymous):

The answer, that is.

OpenStudy (anonymous):

so now I have -3

OpenStudy (anonymous):

\[4x+2y=-1~\iff~4x+2(1-2x)=-1\] \[4x+2-4x=-1\\ 2=-1\] Is that true?

OpenStudy (phi):

what did you get for 2(1-2x) part ?

OpenStudy (anonymous):

I got 2-4x ?

OpenStudy (phi):

and 4x+2 -4x is ?

OpenStudy (anonymous):

x+2 ?

OpenStudy (phi):

4 x's take away 4 x's leaves how many x's ?

OpenStudy (anonymous):

Ohhh 0 X's

OpenStudy (phi):

and 4x+2 -4x is ?

OpenStudy (anonymous):

2 ?

OpenStudy (phi):

yes, so you started with 4x+2(1-2x)= -1 and simplified the left side to 2 2 = -1 that is not true... it is a sign that there is *no solution*

OpenStudy (anonymous):

so the answer is There is no x value as there is no solution ?

OpenStudy (phi):

exactly

OpenStudy (anonymous):

thank you !